Zner062220:

A man deposits P650 every end of 6 months in an account paying 5 1/2 % interest compounded semiannually. What amount is in the account at the end of 5 years and 6 months?

Imagine:

Formula you're going to want to use; \[A=P(1+\frac{ r }{ n })^{nt}\]

Imagine:

\[P=650\] So install "P" into the equation; \[A=650(1+\frac{ r }{ n })^{nt}\] Then Install your interest rate; \[A=650(1+\frac{ 5~1/2 percent }{ ? })^{nt}\] Then Install the Time interest applied per month; \[A=650(1+\frac{ 5~1/2 }{ 0.91 })^{(0.91)(t)}\] Then install the time-period of elapsed time; \[A=650(1+\frac{ 5~1/2 }{ 0.91 })^{(0.91)(5)}\] Use the numbers according to get your final answer.

darkknight:

1 sec, where did .91 come from? could you explain I can't quite follow

Imagine:

dividing how much it gets per month, in a matter of 6 months, it goes up 5 1/2 percent. So, divide it into 6 to get your monthly average.

darkknight:

You don't know what n is then, its number of times compounded per year or month you choose,

imagine wrote:
dividing how much it gets per month, in a matter of 6 months, it goes up 5 1/2 percent. So, divide it into 6 to get your monthly average.
yeah not sure where you got that from. Your P is correct Interest rate should be expressed in decimal, divide 5 1/2 by 100 because it is incorrect the way it is atm. You want to convert time into months or years, depends on what you want to do just know n is number of times interest is compounded per unit time so 6 if you will do monthly or 1/2 if yearly Plug in the right number for time, its not 5

darkknight:

imagine wrote:
\[P=650\] So install "P" into the equation; \[A=650(1+\frac{ r }{ n })^{nt}\] Then Install your interest rate; \[A=650(1+\frac{ 5~1/2 percent }{ ? })^{nt}\] Then Install the Time interest applied per month; \[A=650(1+\frac{ 5~1/2 }{ 0.91 })^{(0.91)(t)}\] Then install the time-period of elapsed time; \[A=650(1+\frac{ 5~1/2 }{ 0.91 })^{(0.91)(5)}\] Use the numbers according to get your final answer.
This answer is wrong in case anyone gets confused

Morahkemz:

Hi, the question requires you to find the future value of an ordinary annuity. It's an ordinary annuity because it's a series of payments made at the end of each period. FV=pmt[((1+r)^n)-1]/r Fv is the future value which we are looking for Pmt is the periodic payment=650 r is the rate per period. Since interest is compounded semi annually, r=0.055/2=0.0275 n is the total number of deposits =11 (2 deposits per year, for 5 years and 6 months) Substituting FV=650*[((1+0.0275)^11)-1]/0.0275 FV=650*[(1.0275^11)-1]/0.0275 FV=650*(1.34772-1)/0.0275 FV=8218.87 The account will have P8218.87 at the end of 5 years and 6 months

surjithayer:

\[amount=650(1.0275)^{11}+650(1.0275)^{10}+...+650(1.0275)^1\] \[=650(1.0275)\frac{ 1.0275^{11}-1 }{ 1.0275-1}\approx 650(1.0275)\frac{ 1.34772143598-1 }{ 0.0275 }\] \[\approx8444.89 \]